# IPAD: Inverse Prompt for AI Detection - A Robust and Explainable LLM-Generated Text Detector

### v1

@@ 4.2 Prompt Inverter techniques and applications
Prompt extraction techniques aim to reverse-engineer the prompts that generate specific outputs from LLMs. Approaches include black-box methods like output2prompt Zhang et al. (2024a), which extracts prompts based on model outputs without access to internal data, and logit-based methods like logit2prompt Mitka (2024), which rely on next-token probabilities but are constrained by access to logits. Adversarial methods can bypass some defenses but are model-specific and fragile Zhang et al. (2024d). Despite the success of some zero-shot LLM-inversion based methods Li and Klabjan (2024); Yu et al. (2024), they are mostly naive usage of prompting LLMs, which makes them poor in prompt extraction accuracy and robustness.
@@ 5 Conclusion
This paper introduces  IPAD (Inverse Prompt for AI Detection), a framework consisting of a  Prompt Inverter that identifies predicted prompts that could have generated the input text, and a  Distinguisher that examines how well the input texts align with the predicted prompts. This design enables explainable evidence chains tracing unavailable in existing black-box detectors. Empirical results show that IPAD surpasses the baselines on all in-distribution, OOD, and attacked data. Furthermore, the  Distinguisher (version2) -  Regeneration Comparator outperforms the  Distinguisher (version1) -  Prompt-Text Consistency Verifier, especially on OOD and attacked data. While the local alignment in veresion1 approach provides explicit interpretability, it is more sensitive to adversarial attacks. In contrast, the global distribution in veresion2 matching approach implicitly learns generative LLM’s distributional properties, which offers more robustness while maintaining explainability. This insight suggests that combining self-consistency checks of generative models with multi-step reasoning for evidential explainability holds promise for future AI detection systems in real-world scenarios. A user study reveals that IPAD enhances trust and transparency by allowing users to examine decision-making evidence. Overall, IPAD establishes a new paradigm for more robust, reliable, and interpretable AI detection systems to combat the misuse of LLMs.
6 Limitations
While IPAD demonstrates SOTA performance, two limitations warrant discussion: (1) The Prompt Inverter may not fully reconstruct prompts containing explicit in-context learning examples (e.g., formatted demonstrations), as it prioritizes semantic alignment over precise syntactic replication. (2) Since IPAD achieves satisfactory OOD performance (12.65% improvement over baselines) by only adopting essay writing datasets for the fine-tuning of Distinguishers, we strategically deferred the exploration of more datasets. We will incorporate a wider and more diverse range of data in future works to explore if it can enhance robustness even further, including: creative/news domains, and triplet data formats (i.e.,  "Can this {predicted prompt} generate the {Input text} using an LLM? One example generated by the predicted prompt is: {regenerated text}")
Acknowledgments
Appendix A AUROC formula
@@
Appendix B Calculation of Summary Statistics
  • $\bullet$
    IPAD with  Regeneration Comparator outperforms the baselines by 9.73% on in-distribution data. As shown in Table 3, RoBERTa-base has the best average F1 score of (92.9% + 92.8% + 91.3% + 78.9%) / 4. In comparison, the average F1 score for IPAD version 2 is (99.85% + 98.5% + 97.6% + 98.86%) / 4, showing an improvement of 9.73%.
  • $\bullet$
    IPAD with  Regeneration Comparator outperforms the baselines by 12.65% on in-distribution data. As shown in Table 4, RoBERTa-base achieves the highest average AUROC score, but since the F1-score is not available for the baseline, we use the AUROC difference to calculate the improvement, which is (95.65% - 83%) = 12.65%.
  • $\bullet$
    IPAD with  Regeneration Comparator outperforms IPAD with  Prompt-Text Consistency Verifier by 0.13% on out-of-distribution (OOD) data. As shown in Table 4, IPAD version 2 has the highest AUROC of 95.65%, while IPAD version 1 has an AUROC of 95.52%, resulting in a 0.13% difference.
  • $\bullet$
    IPAD with  Regeneration Comparator outperforms IPAD with  Prompt-Text Consistency Verifier by 3.78% on attacked data. As shown in Table 3 (rows 3-4) and Table 4 (rows 6-8), IPAD version 2 achieves the best F1 score and AUROC scores. To calculate the overall attacked dataset score, we calculate the F1 scores for Table 4: 94.82%, 95.35%, 95.31% for IPAD version 2, and 83.58%, 88.34%, and 94.70% for IPAD version 1. The average F1 score difference is thus (94.82% + 95.35% + 95.31% - 83.58% - 88.34% - 94.70% + 97.60% + 98.86% - 97.55% - 98.85%) / 5 = 3.78%.
Appendix C DPIC (decouple prompt and intrinsic characteristics) Prompt Extraction Zero-shot Prompts
@@
Appendix D Linguistic Difference Examples
Figure  3 shows examples where HWT and LGT prompts with different sentence complexity. Figure  4 shows the results of analytical thinking level statistics. Figure  5 shows examples of using different personas and different analytical thinking levels.
Figure 3: Sentence Complexity Examples, where  HWT Prompt stands for prompt generated by the Prompt Inverter from HWT, and  LGT Prompt stands for prompt generated by the Prompt Inverter from LGT. The HWT Prompts have longer sentence lengths, more words with more than three syllabus (as shown in bold), and more stop-words (as shown with underline).
Figure 4: Comparison of different analytical thinking levels, with LGT has higher percentage of level 4 and level 5.
Figure 5: Examples that use different persona usage (above), and different analytical thinking levels (below left has level 2, and below right has level 5, they are prompts generated by the same problem statements).
Appendix E User Study
Figure  6 7 and  8 shows the screenshots of online AI detectors. Figure  9 shows the questionnaire questions. Figure  10 shows the user guide.
E.1 Online AI Detectors Screenshots
Figure 6: GPTZero Online Detector Screenshot
Figure 7: Quillbot Online Detector Screenshot
Figure 8: Scribbr Online Detector Screenshot
E.2 Questionnaire questions
Figure 9: Questionnaire questions
Figure 10: User Study User guide

### v2

@@ 4.2 Prompt Inverter techniques and applications
Prompt extraction techniques aim to reverse-engineer the prompts that generate specific outputs from LLMs. Approaches include black-box methods like output2prompt Zhang et al. (2024a), which extracts prompts based on model outputs without access to internal data, and logit-based methods like logit2prompt Mitka (2024), which rely on next-token probabilities but are constrained by access to logits. Adversarial methods can bypass some defenses but are model-specific and fragile Zhang et al. (2024c). Despite the success of some zero-shot LLM-inversion based methods Li and Klabjan (2024); Yu et al. (2024), they are mostly naive usage of prompting LLMs, which makes them poor in prompt extraction accuracy and robustness.
@@ 5 Conclusion
This paper introduces  IPAD (Inverse Prompt for AI Detection), a framework consisting of a  Prompt Inverter that identifies predicted prompts that could have generated the input text, and two  Distinguishers that examines how well the input texts align with the predicted prompts. One is the  Prompt-Text Consistency Verifier (PTCV) which evaluates direct alignment between predicted prompts and input text, and the other is  Regeneration Comparator (RC) that examines content similarity by comparing input texts with the corresponding regenerated texts. Empirical evaluations demonstrate that IPAD outperforms the strongest baselines by 9.05% (Average Recall) on in-distribution data, 12.93% (AUROC) on out-of-distribution (OOD) data, and 5.48% (AUROC) on attacked data. The combination of the two modules suggests that combining self-consistency checks of generative models with multi-step reasoning for evidential explainability holds promise for future AI detection systems in real-world scenarios. An interpretability assessment reveals that IPAD enhances trust and transparency by allowing users to examine decision-making evidence.
Limitations
While IPAD demonstrates SOTA performance, two limitations warrant discussion: (1) The Prompt Inverter may not fully reconstruct prompts containing explicit in-context learning examples, as it prioritizes semantic alignment over precise syntactic replication. (2) While IPAD achieves strong performance across diverse datasets, it relies on LLMs, making it more computationally expensive compared to lightweight detectors such as RoBERTa or HC3. However, compared other detectors compared with LLMs, such as DPIC, IPAD is more lightweight since it calls the open-sources light-weight Phi-3 model.
Appendix A Fine-tune Dataset
Prompt Inverter Dataset.
We use the following four datasets, with the first three datasets enhance the model’s generalization to recover the prompts, while the last dataset improves performance on essay-related tasks.
  • Instructions-2M Morris et al. (2024), a collection of 2 million user prompts and system prompts, from which we used 30,000 prompts.
  • ShareGPT Zhang et al. (2024d), an open platform where users share ChatGPT prompts and responses, from which we used 500 samples.
  • Unnatural Instructions Zhang et al. (2024d), a dataset of creative instructions generated by OpenAI’s models, from which we used 500 samples.
The first three datasets aims to enhance the general querying capability of the Prompt Inverter, and are all released under the MIT license. All the samples we used are the same to the samples randomly selected in  Zhang et al. (2024a). The last dataset aims to enhance the familiarity of the Prompt Inverter with the data of the essay to detect the LLM-generated essays, and are created and examined by  Koike et al. (2024), We specifically used the LLM-generated essays and problem statements for this supervised fine-tuning (SFT). There are 45,400 training pairs in total.
Given that essay data are diverse, we utilize only the OUTFOX dataset Koike et al. (2024). To adapt this dataset for training our Distinguisher, we enhance it to align with the model’s requirements. The original dataset consists of 14,400 training triplets of essay problem statements, student-written essays, and LLM-generated essays. To further process the data, we apply the Prompt Inverter to both student-written and LLM-generated essays, generating corresponding Predicted Prompts. These Predicted Prompts are then used to regenerate texts via ChatGPT, i.e. gpt-3.5-turbo. Following this procedure, we construct a total of 28,800 training samples, with an equal distribution of positive and negative examples (14,400 each).
The final dataset is structured as follows:
Table 7: Instruction, input/output structure, and inference outputs of each fine-tuned module. $T$ is the input text, $P$ the predicted prompt, and $T^{\prime}$ the regenerated text.
Field Prompt Inverter PTCV RC
Instruction "What is the prompt $P$ that generates the Input Text $T$?" "Can LLM generate the input text $T$ through the prompt $P$?" "$T^{\prime}$ is generated by LLM, determine whether $T$ is also generated by LLM with a similar prompt."
Input $T$ $(P,T)$ $(T^{\prime},T)$
Output $P$yes"/"no" "yes"/"no"
Output in Inference $P$ $p_{\text{PTCV}}$ $p_{\text{RC}}$
Appendix B Complete Algorithm
The complete inference pipeline is summarized in Algorithm 1.
Algorithm 1 IPAD Detection Procedure
1:Input text $T$; trained modules $\mathbf{f}_{\text{inv}},\mathbf{f}_{\text{PTCV}},\mathbf{f}_{\text{RC}}$; LLM $\mathbf{f}_{\text{LLM}}$; fusion weight $w\in[0,1]$; threshold $\tau\in[0,1]$
2:Prediction $\hat{Y}\in\{\mathrm{HWT},\mathrm{LGT}\}$ and evidence $\mathcal{E}$
3:$P\leftarrow\mathbf{f}_{\text{inv}}(T)$triangleright Inverse-prompt prediction
4:$T^{\prime}\leftarrow\mathbf{f}_{\text{LLM}}(P)$triangleright Regenerate text using P
5:$\boldsymbol{z}^{\text{PTCV}}\leftarrow\mathbf{f}_{\text{PTCV}}(P,T)$
6:$p_{\text{PTCV}}\leftarrow\dfrac{\sigma(\boldsymbol{z}^{\text{PTCV}}_{\text{yes}})}{\sigma(\boldsymbol{z}^{\text{PTCV}}_{\text{yes}})+\sigma(\boldsymbol{z}^{\text{PTCV}}_{\text{no}})}$
7:$\boldsymbol{z}^{\text{RC}}\leftarrow\mathbf{f}_{\text{RC}}(T^{\prime},T)$
8:$p_{\text{RC}}\leftarrow\dfrac{\sigma(\boldsymbol{z}^{\text{RC}}_{\text{yes}})}{\sigma(\boldsymbol{z}^{\text{RC}}_{\text{yes}})+\sigma(\boldsymbol{z}^{\text{RC}}_{\text{no}})}$
9:$\hat{p}\leftarrow w\cdot p_{\text{PTCV}}+(1-w)\cdot p_{\text{RC}}$
10:if $\hat{p}>\tau$ then
11:     $\hat{Y}\leftarrow\mathrm{LGT}$
12:else
13:     $\hat{Y}\leftarrow\mathrm{HWT}$
14:end if
15:$\mathcal{E}\leftarrow\bigl{(}P,\;p_{\text{PTCV}},\;p_{\text{RC}},\;\hat{p}\bigr{)}$
16:return $(\hat{Y},\;\mathcal{E})$
Appendix C AUROC formula
@@
Appendix D Ablation study data structures
Input-only fine-tuning data instructions.
"Is this text generated by LLM?"
Prompt Only fine-tuning data instructions.
"Prompt Inverter predicts prompt that could have generated the input texts. Is this prompt predicted by an input texts written by LLM?"
Ablation Prompt.
"Text A is generated by an LLM. Determine whether Text B is also generated by an LLM using a similar prompt. Meanwhile, determine whether Text B could have been generated from Prompt C using an LLM. Answer with YES or NO."
Appendix E DPIC (decouple prompt and intrinsic characteristics) Prompt Extraction Zero-shot Prompts Yu et al. (2024)
@@ Appendix D Ablation study data structures
Appendix F Comparison with DPIC
Since DPIC has not released its code, data, or models, we are unable to independently evaluate the performance of its classifier. Consequently, we rely on the reported results in the DPIC paper and construct a comparable dataset following their described settings to enable a fair comparison with IPAD. However, due to these limitations, we are unable to apply DPIC to additional datasets for broader evaluation.
To assess the generalization of IPAD, we reconstruct the following datasets, each containing 200 randomly sampled examples: XSum, WritingPrompts, and PubMedQA. For each dataset, we generate texts using three large language models: ChatGPT (gpt-3.5-turbo), GPT-4 (gpt-4), and Claude 3 (claude-3-opus-20240229). Furthermore, the XSum datasets generated by these three models are augmented using two attack methods—DIPPER and Back-Translation—resulting in a total of 15 evaluation datasets.
Table 8: AUROC comparison across tasks (XSum, Writing, PubMed) for ChatGPT, GPT-4, and Claude 3 using various prompt extraction methods.
Method ChatGPT GPT-4 Claude 3
XSum Writing PubMed Avg. XSum Writing PubMed Avg. XSum Writing PubMed Avg.
DPIC (ChatGPT) 1.0000 0.9821 0.9092 0.9634 0.9996 0.9768 0.9438 0.9734 1.0000 0.9950 0.9686 0.9878
DPIC (Vicuna-7B) 0.9976 0.9708 0.8990 0.9558 0.9986 0.9644 0.9394 0.9674 0.9992 0.9943 0.9690 0.9875
IPAD (Version 1) 0.9850 0.9800 0.9250 0.9633 1.0000 0.9700 0.9700 0.9800 1.0000 0.9800 0.9750 0.9850
IPAD (Version 2) 1.0000 0.9850 0.9800 0.9883 1.0000 0.9800 0.9500 0.9767 1.0000 0.9950 1.0000 1.0000
Table 9: AUROC comparison under generation perturbation settings (DIPPER, Back-translation) for each model.
Method ChatGPT GPT-4 Claude 3
Ori. DIPPER Back-trans. Ori. DIPPER Back-trans. Ori. DIPPER Back-trans.
DPIC (ChatGPT) 1.0000 1.0000 0.9972 0.9996 0.9991 0.9931 1.0000 0.9996 0.9878
DPIC (Vicuna-7B) 0.9976 0.9980 0.9889 0.9986 0.9969 0.9903 0.9992 0.9996 0.9979
IPAD (Version 1) 0.9850 0.8900 0.9850 1.0000 0.8950 0.9900 1.0000 0.9250 0.9950
IPAD (Version 2) 1.0000 0.9750 0.9950 0.9800 0.9750 0.9950 1.0000 1.0000 1.0000
Based on the experimental results, IPAD performs well and exhibits notable resistance to adversarial attacks.
IPAD open-sourced all the fine-tuned models, including the Prompt Inverter, and the two versions of distinguishers. Therefore, all the experiment results can be validated and reproduced.
Appendix G Different Linguistic Features of HWT prompts and LGT prompts
One of the primary distinctions between the HWT prompts and the LGT prompts is sentence complexity. LGT prompts are typically more complex, characterized by longer sentence lengths (mean value of 1.514 and 1.794), higher syllable counts (mean values of total syllabus three are 1.572 and 3.042), and more stop-words (mean values of 9.88 and 10.045). HWT prompts, on the other hand, are characterized by shorter, less complex sentences that are easier to process and understand.
Beyond the differences in syntactic complexity, we also explored variables in LIWC. We did the difference comparison by using HWT and LGT prompts as inputs for ChatGPT, for example, instructing with the prompts ’determine the pronoun usage of this sentence, answer first person, second person, or third person’ and ’determine the level of analytical thinking of these sentences, answer a number from 1 to 5’. The results show that there are distinguish difference in pronoun usage and analytical thinking level. The HWT prompts frequently use second-person pronouns (e.g., ’you’) - 75 occurrences per 1,000 prompts - due to the subjective tone often employed in HWT. In contrast, LGT prompts primarily feature first- and third-person pronouns, with second-person pronouns appearing only 2 per 1,000 prompts. LGT prompts typically present instructions and questions in a more objective manner. LGT prompts show higher analytical thinking levels than HWT prompts. With level 1 as the lowest and level 5 as the highest, LGT has 68.9% of level 4 and 24.3% of level 5, but HWT has only 48.0% of level 4, and 0.8% of level 5. It suggests that LGT prompts encourage more analytical thinking, while HWT prompts tend to focus more on concrete examples, with less emphasis on critical analysis.
Appendix H IPAD and DPIC prompt inverter examples
Table 10: IPAD and DPIC prompt inverter examples
Input IPAD DPIC
The IPC opened proceedings against the National Paralympic Committee of Russia after a report claimed the country had operated a widespread doping programme.A decision on any ban will come in the week commencing 1 August.The International Olympic Committee (IOC) has opted against a blanket ban."I can assure you that our board will take the right decision in the interest of sport and the interest of the Paralympic movement moving forward," said Craven.On Sunday, the IOC said it would leave it up to the governing bodies of individual sports to decide if Russian competitors are clean and should be allowed to take part.But Craven, himself a member of the IOC, was critical of that decision and said the IPC would not necessarily follow suit."I am disappointed in their decision, but that is a personal view," he added."We have to acknowledge their right to take such a decision. This is ultra-serious. I don’t think there has been a situation in the past where you have had institutional doping on such a scale."We believe the Russian NPC is either unwilling or unable to uphold the IPC anti-doping code, which is in line with the World Anti-Doping Agency code, so that is what they have to respond to."Canadian law professor Richard McLaren’s report, published last week, claimed Russia operated a state-sponsored doping programme from 2011 to 2015.The IPC said it acted after McLaren provided the names of the athletes associated with the 35 "disappearing positive samples" from the Moscow laboratory highlighted in the report.Nineteen samples potentially doctored as part of the sample-swapping regime during the 2014 Sochi Paralympic Winter Games have been sent for further analysis.Russia will have up to 21 days to appeal against any IPC decision, with the Rio Paralympics due to begin on 7 September. Explain the actions taken by the International Paralympic Committee against the Russian Paralympic Committee and the reasons behind these actions. What were the key findings of Richard McLaren’s report regarding Russia’s doping programme?
The world came crashing down in minutes. Many of us were asleep when it happened, and did n’ t find out about it until later. When we awoke, we saw the carnage spread through the land, and we wept. There were n’ t many of us left, but what few there were managed to find each other over the Internet. We gathered together in what remained of a major city on the East Coast of what was once the United States. It took us time, but we eventually began to rebuild. The brightest among those who survived thought to ask " Why, " while most of us were content with just surviving. Years passed, and no link was found between us. Eventually, those who had the question resigned themselves to the fact that they would never know. They went to their new homes, and tried to integrate themselves as best they could into the new society. It was n’ t until 14 years after the event happened that the connection was discovered, quite by accident. One of the former questioners had taken a job as a mover, and was helping a fellow survivor move into a newly cleaned house. Sticking out of one of the boxes was the missing link. " Oh, you used to shop at Bad Dragon too? " Moments later, the realization struck him. Write an essay about a time when you and your friends were the only survivors of a catastrophic event that wiped out most of the world’s population. Describe how you and your friends coped with the aftermath and the challenges you faced in rebuilding society. What was the event that caused the world to come crashing down?
In an alternate timeline, a second sentient race evolved in parallel with humans. These beings, known as the Avralians, possessed extraordinary abilities and resided in the hidden corners of the Earth. For centuries, unaware of each other’s existence, humans and Avralians progressed independently. However, fate intervened one fateful day when an Avralian girl named Elara stumbled upon a group of human explorers deep in the forest. With wide-eyed wonder, she approached them cautiously. The humans, initially startled, soon recognized Elara’s unique nature. They befriended her, and through her, the two races began to unravel the secrets of their intertwined existence. As time passed, harmony emerged between the two races, and mutual respect solidified their bond. The Avralians shared their ancient knowledge and wisdom, while humans brought innovation and technology. Together, they worked towards a world where unity triumphed over differences. However, not all were convinced of this newfound alliance. A faction within the human population, fueled by fear and mistrust, sought to exploit Avralian abilities for personal gain. Conflict brewed, threatening the fragile equilibrium. Elara, fueled by her pure heart and belief in unity, emerged as a beacon of hope. With a small band of supporters, she embarked on a journey to bridge the gap and foster understanding. Battles were fought, sacrifices made, but ultimately, Elara’s message prevailed. Humans and Avralians learned to cherish their diversity and forge a future marked by collaboration and empathy. The world transformed into a tapestry of coexistence, where magnificent cities stood as testaments to unity and cultural exchange. Humans and Avralians moved freely through bustling markets, sharing knowledge, stories, and laughter. Together, they faced global challenges, from climate crises to epidemics, with unwavering determination. Write an essay describing an alternate timeline in which a second sentient race evolved in parallel with humans, exploring the potential interactions and conflicts between the two species. How did Elara manage to convince both races to embrace unity despite the conflict?
Both times I had the banana pepper appetizer, which is great and goes really well with the FRESH and delicious bread and cheese they give you at the start of your meal. nnFor entrees, me and my girlfriend have had mixed experience. I’ve had the fish sandwich (very good) and the eggplant parm sandwich (okay). My girlfriend got the salad with bread and basil on it, but the basil was over powering and the bread was soggy with the dressing. nnThe service is also a mixed bag. The first time our server went out of her way to take care of us and even MADE me cocktail sauce for my fish sandwich. The second time, the server was lackluster, didn’t know anything about the menu and wasn’t able to take proper care of us. nnI would return to Papa J’s, but I my terrible experience last time isn’t enough to say it would be my first pick of places to eat around Carnegie/Robinson. This was a great place to stop for a quick lunch. The lines were not too long for the sandwiches they had and they had a wide selection of bagels if you wanted a bagel sandwich. With a great front patio for enjoying your food, it was a relaxing place to stop. Write a review for it. What made the banana pepper appetizer stand out to you compared to other starters?
Abstract: This article explores the longstanding debate between Einstein’s theory of general relativity and Maxwell’s theory of electromagnetism regarding the nature of gravitation. The central question addressed is whether gravitation is best understood as a curvature of space, a field in flat space, or perhaps a combination of both concepts. Drawing upon a comprehensive analysis of the theoretical framework and empirical evidence, the article presents a nuanced examination of the arguments put forth by Einstein and Maxwell.The article begins by discussing Einstein’s general theory of relativity, which proposes that gravitation arises from the curvature of spacetime caused by mass and energy. It outlines the mathematical formalism used to describe this curvature and highlights the key predictions and experimental confirmations of the theory. Conversely, the article delves into Maxwell’s electromagnetic theory, which suggests that gravitation may be explained as a fundamental force mediated by a field propagating through flat space, similar to electromagnetic fields.Further, the article explores the distinctive features and limitations of each theory. It scrutinizes the conceptual foundations, mathematical rigor, and empirical support for both approaches, highlighting their respective strengths and weaknesses. Moreover, the article examines attempts to reconcile the two theories into a unified framework, such as the development of theories of quantum gravity.By critically evaluating the arguments and evidence from both camps, this article aims to offer a comprehensive assessment of the question regarding the nature of gravitation. Based on the analysis presented, it becomes evident that both Einstein’s theory of general relativity and Maxwell’s theory of electromagnetism provide valuable insights into the phenomenon of gravitation. Write a paper abstract to explain the debate between Einstein’s theory of general relativity and Maxwell’s theory of electromagnetism regarding the nature of gravitation, and argue for which theory is more likely to be correct based on the evidence presented in the essay statement. What are the main challenges in reconciling Einstein’s theory of general relativity with Maxwell’s theory of electromagnetism in explaining gravitation?